Following are the demand and supply function of a market


QDA = 8000 - 1000Px, QSA= - 4000+2000Px


i) Find out the market clearing price and quantity.


ii) Plot, on one set of axes, the market demand curve and the market supply curve for commodity A and show the equilibrium point.


iii) Is the equilibrium stable? Explain.


1
Expert's answer
2021-11-14T17:38:15-0500

(I)

"Qd=8000-1000p\\Qs=-4000+2000p\\Qd=Qs\\8000-1000p=-4000+2000p\\8000+4000=2000p+1000p\\12000=3000p\\p =4\\Q=8000-(1000×4)\\=4000units"

(ii)


(iii) Equilibrium is stable. This is because at point E demand and supply are at balance.